Hi. My name is Lindsey and I am trying to figure out how to solve this.
3(x-6)+8=5(x+2)-8 so far i have this.
3x-18+8=5x+10-8

Found 2 solutions by algebrahouse.com, stanbon:
Answer by algebrahouse.com(1659)

3(x - 6)+ 8 = 5(x + 2) - 8
3x - 18 + 8 = 5x + 10 - 8 {used distributive property}
3x - 10 = 5x + 2 {combined like terms}
-2x = 12 {subtracted 5x and added 10 to each side}
x = -6 {divided each side by -2}

For more help from me, visit:
www.algebrahouse.com


Answer by stanbon(75887)

3x-18+8=5x+10-8
-----
3x-10 = 5x+2
---
2x = -12
x = -6
============
Cheers,
Stan H.
============